What is the formula to solve this?

The Perimeter of a rectangle is 600. The length of the rectangle is 2 more than 3 times the width.

Answer by venugopalramana(3286)

What is the formula to solve this?
The Perimeter of a rectangle is 600. The length of the rectangle is 2 more than 3 times the width.
PERIMETER=2*(L+W)
LET WIDTH =W
3 TIMES WIDTH =3W
2 MORE THAN THAT =3W+2=L
SO PERIMETER =2(3W+2+W)=8W+4=600
8W=600-4=596
W=596/8=74.5
L=3*74.5+2=223.5+2=225.5